OpenStudy (vjm16):

15 times the square of a non zero number is equal to 20 times the number

OpenStudy (legomyego180):

\[15x^2=20x\]

OpenStudy (legomyego180):

Solve for x over the real numbers: 15 x^2 = 20 x Subtract 20 x from both sides: 15 x^2-20 x = 0 Factor x and constant terms from the left hand side: 5 x (3 x-4) = 0 Divide both sides by 5: x (3 x-4) = 0 Split into two equations: x = 0 or 3 x-4 = 0 Add 4 to both sides: x = 0 or 3 x = 4 Divide both sides by 3: Answer: | | x = 0 or x = 4/3